OpenStudy (anonymous):

I need help with a final and I have to get these problems down... Can someone take the time to help me with these ..???? 4(x-7)=7x-12 distribute. I need knowledge and what these problems mean so I can solve or set up as asked.

OpenStudy (anonymous):

mul;tiply the number in front of the parentheses times each term inside 4x-28=7x-12 get the variables (x) to one side and constants to the other 3x=-16 x=-16/3=-5 1/3

OpenStudy (anonymous):

4(x-7)=7x-12 4x-4*7=7x-12 4x-28=7x-12 -28+12=7x-4x -16=3x -15/3 = x

OpenStudy (anonymous):

u guys arre smarty pants. lol

OpenStudy (anonymous):

Lol. This is elementary maths!

OpenStudy (anonymous):

haha. the truth is that i didnt even read the question.

OpenStudy (anonymous):

Lol.

OpenStudy (anonymous):

haha